OpenStudy (anonymous):

Find the range: y = -x2 - 2x + 5

OpenStudy (anonymous):

\[y=-x^2-2x+5\] is a parabola that opens down you need the second coordinate of the vertex the first coordinate is \(-\frac{b}{2a}\) which in your case is \[-\frac{-2}{2\times (-1)}=-1\]

OpenStudy (anonymous):

the second coordinate of the vertex is what you get for \(y\) when you replace \(x\) by \(-1\)

OpenStudy (anonymous):

\[y = -x^2 - 2x + 5\]\[y=(-1)^2-2\times (-1)+5\] is what you need to compute

OpenStudy (anonymous):

oops that was not right, compute \[y=-(-1)^2-2\times (-1)+5\]

OpenStudy (anonymous):

should get \[-1+2+5=6\]